OpenStudy (anonymous):

Electrical Current. The current I in an electrical conductor varies inversely as the resistance R of the conductor. The current is ¼ ampere when the resistance is 3840 ohms. What is the current when the resistance is 2160 ohms?

OpenStudy (anonymous):

In first case, I1 = 1/4 A = 0.25 A R1 = 3840 V = ?? V = R1*I1 = 3840 * 0.25 = 960V In second case V = 960V R2 = 2160 I2 = ?? I2 = V/R2 = 960/2160 = 0.44 A